Session: Feb 2003 (session 1)
Wavelength: 1.3 cm
To give feedback on this experiment, click
Onsala (M. Lindqvist at Sun Feb 9 08:01:00 2003) Success :
No known problems.
Noto (Pietro Cassaro at Mon Feb 10 11:32:05 2003) Success :
Jodrell_Bank (Alastair Gunn at Mon Feb 10 14:48:53 2003) Minor failures :
Schedule was stopped and restarted at 19:09 to correct impossible
"alt,alt" IFD setting. This was reset to "nor,nor", LOs reset and
probably means the polarisations are swapped. PI seems to have used
a non-standard (and wrong) setup. Hence, data probably only good
from 19:09 restart. No other known problems.
Medicina (g.Maccaferri at Mon Feb 10 14:56:33 2003) Minor failures : FS
Computer problem at 23:48. Schedule interrupted half an hour before the end.
Robledo (Z.P. (for Esther Moll) at Wed Feb 12 14:52:43 2003) Minor failures : LATE SLEW
1) During countdown RAC60 crashed, control of K-band
microwaves, MMS and RF Matrix Switch was not possible.
The MMS had to be configured at the k-band cone and the RF Matrix
Switch had to be bypassed. At 21:44Z RAC60 computer was finally recovered.
At 22:19Z (after some MMS and Post-amplifier attenuation
adjustments) the system was recovered. (see DR# M101376).
2) Level on video converter one was stable since
22:19Z, when configuration was finally completed
3) Diode control worked for the whole activity
4) Full_cal procedure worked (was performed at the
end of the experiment)
5) The formatter was in sync
6) Tape positioning was correct
7) For scans 1133 and 1175 late on source events
happened and for scans 1151 and 1165 sources were completely lost. Reason:
not enough time was available on schedule file for arriving on
time to those sources. (See DR#M101387)
8) Source for scan 1295 was lost due to the
communication problem between EAC and DSS63 antenna.
9) For scans 1606 and 1638 late on source events
happened, and for scan 1624 source was completely lost. Reason: not enough
time was available on schedule file for arriving on time to
those sources. (See DR#M101387)
> Weather condition: Clear
> Station Time delay: -7.807
> Recorder used: #1
> K-Band (LCP) System Temp:128.3
> K-Band (LCP) Y-factor:3.9
> DR #:M101376
> RAC60 crashed during countdown. See DR#M101376
> Observation stopped at 01:04 per agreement with
> Principal Investigator (antenna pointing to mountain).